OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $395. How many sophomores bought tickets? 55 75 90 95

OpenStudy (nurali):

f + s =150 f = 150 - s ----------------- eq 1 2f + 3s = 395 subst eq 1 2(150 - s) + 3s = 395 300 - 2s + 3s = 395 s = 395 - 300 s = 95